### v2

# EvRWKV: A Continuous Interactive RWKV Framework for Effective Event-Guided Low-Light Image Enhancement

Abstract
Capturing high-quality visual content under low-light conditions remains a challenging problem due to severe noise and underexposure, which degrade the performance of downstream applications. Traditional frame-based low-light image enhancement methods often amplify noise or fail to preserve structural details. Event cameras, offering high dynamic range and microsecond temporal resolution by asynchronously capturing brightness changes, emerge as a promising complement for low-light imaging. However, existing fusion methods fail to fully exploit this synergy, either by forcing modalities into a shared representation too early or by losing vital low-level correlations through isolated processing. To address these challenges, we propose EvRWKV, a novel framework that enables continuous cross-modal interaction through dual-domain processing. Our approach incorporates a Cross-RWKV module, leveraging the Receptance Weighted Key Value (RWKV) architecture for fine-grained temporal and cross-modal fusion, and an Event Image Spectral Fusion Enhancer (EISFE) module, which jointly performs adaptive frequency-domain noise suppression and spatial-domain deformable convolution alignment. This continuous interaction maintains feature consistency from low-level textures to high-level semantics. Extensive qualitative and quantitative evaluations on real-world low-light datasets (SDE, SDSD, RELED) demonstrate that EvRWKV achieves state-of-the-art performance, effectively enhancing image quality by suppressing noise, restoring structural details, and improving visual clarity in challenging low-light conditions.
Index Terms:
Low-light image enhancement, event cameras, cross-modal fusion, RWKV, dual-domain processing.
Figure 1: Overall framework of event-guided LLIE approaches. (a) Early fusion: combining image and event data at the input. (b) Late fusion: processing image and event data separately and merging features. (c) Continuous interaction fusion: enabling ongoing interaction between image and event data.
I Introduction
Capturing high-quality visual content under low-light conditions remains a critical yet challenging task in computer vision [26]. Images and videos acquired in such environments often suffer from severe noise, motion blur, and underexposure [26], which degrade downstream applications such as autonomous driving and surveillance. Traditional frame-based low-light image enhancement (LLIE) methods [23, 47, 9, 50] have made progress, but often struggle to effectively address the inherent complexities of real-world low-light scenarios. Many methods [55, 3, 49, 33, 27]can amplify existing noise while attempting to brighten the image, leading to visually unappealing results. Furthermore, approaches relying on synthetic data for training often fail to generalize well to the diverse and unpredictable nature of real-world low-light scenes.
Event cameras, using bio-inspired sensors [32], are widely used in the field of computer vision due to their high dynamic range and microsecond level time resolution [43, 59], such as image restoration [17], video super-resolution [37], optical flow [38], deblurring [18, 19, 6], etc., and have also become a promising alternative to low-light imaging scenes. By asynchronously capturing per-pixel brightness changes, events excel at preserving structural details even in extreme darkness and are less susceptible to motion blur. This makes them particularly attractive for low-light image enhancement, where traditional cameras struggle due to signal-to-noise ratio limitations and motion artifacts. However, exploiting event cameras for LLIE is not straightforward. Existing methods [44, 42] that rely solely on event cameras often face limitations due to the sparse nature of the event data, which can lead to poor spatial resolution and difficulty in handling noise under extreme low light conditions. Furthermore, fusion techniques combining events and images often use simplistic approaches [17, 46, 35, 30], which fail to effectively capture the complex spatial-temporal interactions between them. As a result, these methods do not fully leverage the complementary strengths of event and image data, limiting their performance in LLIE tasks.
Figure 2: A challenging example from our dataset containing an extremely low-light image (a) and sparse events (b). Compared with the result from the SOTA event-guided method EvLight (c), our EvRWKV (d) not only recovers the wheel details in the dark areas (e.g., the wheel) but also preserves edge details (e.g., the white line on the floor).
We argue that a core challenge lies in designing a fusion architecture that respects the profound heterogeneity of event and image modalities while enabling meaningful synergy between them. As shown in Fig. 1 (b), prevailing methods often adopt a decoupled design [30, 20, 45, 36, 17], processing event and image streams through separate, parallel pathways. This architectural choice inherently inhibits continuous interaction, restricting feature exchange to a single terminal fusion step. Consequently, vital low-level complementary correlations are lost during independent hierarchical processing. These include transient event edges corresponding to faint image textures. By the time the features are combined, they are too abstract to benefit from this lost complementary detail, leading to underutilized complementarity, semantic misalignment, and suboptimal enhancement, as shown in Fig. 2 (c).
Conversely, as shown in Fig. 1 (a), strategies [31, 35] that merge modalities at the input level struggle with a fundamental mismatch in data structure. Events are temporally dense but spatially sparse [32], whereas images offer dense spatial information, but are temporally discrete [26]. Forcing these inherently complementary representations into a common format at the outset creates a representational bottleneck, obscuring their distinct characteristics and preventing networks from effectively learning complementary relationships. This approach may also amplify noise from both sources [10, 26], degrading the fused representation before feature extraction begins. The fundamental limitation across these paradigms is the absence of a continuous adaptive interaction between the modalities throughout the feature learning process.
To address this challenge, we introduce EvRWKV, a novel framework engineered for continuous cross-modal interaction through dual-domain processing. Instead of treating fusion as a single-stage process, EvRWKV establishes persistent cross-modal interactions between the RGB data and event streams at every network layer. This ensures that features from each modality dynamically inform and refine one another throughout the entire representation learning hierarchy. At the core of our approach, the Cross-RWKV module leverages the Receptance Weighted Key Value (RWKV) architecture [40] to enable fine-grained interaction between event and image features, capturing temporal dependencies and ensuring consistent spatiotemporal alignment. Complementing this, the Event Image Spectral Fusion Enhancer (EISFE) module addresses noise and structural degradation by applying an Adaptive Gaussian Filter in the frequency domain to suppress noise while preserving key structural details, and using deformable convolutions with learnable offset fields in the spatial domain to align and fuse high-frequency details from both modalities. By ensuring dynamic interaction across all stages of feature extraction—from low-level edges and textures, which are calibrated in real time, to mid-level structures that compensate for each other’s deficiencies, and high-level abstract features that are optimized for consistency—EvRWKV achieves seamless cross-modal fusion. This end-to-end collaboration maintains continuous feature alignment, preventing early information loss or late-stage disconnection, thus delivering robust LLIE that effectively preserves fine details, ensures global consistency, and suppresses noise and structural degradation.
In summary, EvRWKV represents a significant advancement in LLIE by effectively harnessing the synergistic potential of event and image cameras through a unified, cross-modal, and cross-domain architecture. The contributions of this work are as follows:
  • We propose EvRWKV, a novel framework that establishes continuous cross-modal interaction through dual-domain processing for LLIE, establishing intrinsic information exchange between event and image modalities.
  • We present Cross-RWKV, an RWKV-based backbone that captures temporal dynamics and performs cross-modal fusion through simultaneous event-image processing for enhanced low-light performance.
  • We develop EISFE, a module operating in frequency and spatial domains to achieve noise mitigation, structure enhancement, and multi-modal feature alignment.
  • We demonstrate that EvRWKV achieves state-of-the-art (SOTA) performance on LLIE across real-world datasets (SDE, SDSD, RELED), effectively suppressing noise and improving visual quality.
II Related Work
II-A Frame-based LLIE
LLIE aims to improve the visibility and quality of images captured in dimly lit environments. Traditional frame-based LLIE methods primarily operate on standard image frames. These methods can be broadly categorized into histogram equalization [2, 1, 4], Gama Correction [16, 41] and Retinex-based algorithms [12, 54, 9, 58, 33]. The first two methods directly enhance the intensity and contrast of low light image, while the Retinex-based algorithm models the image as a combination of illumination and reflectivity. Although effective in improving visibility, these approaches often encounter limitations in handling complex lighting conditions, especially in scenes with varying illumination or dark regions. Moreover, these methods may amplify noise, particularly in low-light regions, leading to undesirable artifacts.
Recent advancements have been propelled by deep learning, particularly with convolutional neural networks (CNNs) [9, 28, 5] and Transformer architectures [55, 3, 60], which have led to the development of more sophisticated frame-based LLIE methods. Chen et al. [5] established a low light image dataset and proposed a full convolution network for enhancement. Wang et al. [48] proposed to enhance underexposed photos by learning the illumination map. Wang et al. [47] collect SDSD dataset by using mechatronic system and proposed a framework integrating progressive alignment and Retinex-based illumination prediction. Cai et al. [3] propose a one-stage Transformer framework Retinexformer for LLIE, leveraging Retinex theory and illumination-guided attention to suppress noise. However, these frame-only methods fundamentally suffer from information loss in dark regions due to sensor noise floor limitations, leading to irreversible texture degradation and motion blur amplification.
II-B Event-guided LLIE
The emergence of event cameras has significantly revolutionized low-light perception by offering a new paradigm in image sensing. Unlike traditional frame-based cameras that capture images at fixed intervals, event cameras detect changes in the scene asynchronously, producing a continuous stream of events that encode pixel-level intensity changes with high temporal resolution. Rebecq et al. [42] propose a recurrent network to reconstruct high speed and high dynamic range videos from event camera data. Event-based approaches for LLE like NER-Net [34] leveraged event streams’ temporal continuity to address non-uniform illumination via learnable timestamp calibration. Subsequent works explored hybrid event-frame fusion strategies. Jiang et al. [17] established joint feature learning through attention, while [30] introduced the SDE dataset with precise spatial-temporal alignment and proposed SNR-guided feature selection in EvLight framework. Kim et al. [20] propose an event-guided end-to-end framework for joint low-light video enhancement and deblurring on the real-world RELED dataset, utilizing temporal alignment and cross-modal spectral filtering. However, current fusion architectures are still limited by methods that rely on basic fusion strategies, which fail to effectively separate modality-specific noise patterns. Additionally, many of these approaches suffer from inflexible fusion operators that lack adaptive cross-modal interaction mechanisms. Furthermore, they often underutilize the event streams’ intrinsic motion information, hindering comprehensive dynamic scene modeling.
II-C Receptance Weighted Key Value
The Receptance Weighted Key Value (RWKV) architecture [40] emerges as a transformative sequential modeling paradigm, synergizing the parallel computation advantages of transformers with the memory efficiency of recurrent networks. Departing from conventional attention mechanisms constrained by quadratic complexity, RWKV introduces time-shifted receptance gates and exponentially decaying key projections to maintain hidden states [39], enabling linear computational scaling with sequence length. In computer vision, RWKV has demonstrated remarkable versatility across diverse tasks [29]. For instance, Restore-RWKV [57] employs an omnidirectional token translation layer and recurrent WKV attention to restore medical images [56] by capturing spatial dependencies, while PointRWKV [13] leverages linear complexity and multi-head matrix states to enhance geometric feature extraction in 3D point cloud processing. Vision-RWKV [8] further extends its capabilities to high-resolution image analysis by reducing spatial aggregation complexity, outperforming traditional Vision Transformers (ViTs). In real-time video analytics, the TLS-RWKV framework [61] achieves efficient temporal pattern recognition for action detection. StyleRWKV [7] utilizes RWKV models for efficient style transfer, enhancing global and local context while maintaining low memory usage and linear time complexity.
Despite these advances, RWKV’s application to cross-modal fusion scenarios—particularly for processing asynchronous, heterogeneous data streams like images and events—remains underdeveloped. Existing RWKV variants struggle with asynchronous temporal alignment and spatial-temporal sparsity gaps. While recent works like RWKV-CLIP [11] explore vision-language alignment, they primarily focus on static multimodal interactions. Our work proposes a novel perspective: establishing continuous event-image synergy, where events resolve temporal ambiguities in blurry regions through high-frequency temporal updates, while images anchor sparse event data to structural priors via spatial-semantic grounding. This approach aligns with RWKV’s inherent strength in modeling long-range dependencies while mitigating its limitations in handling asynchronous modalities, opening new pathways for applications in motion deblurring, LLIE, and neuromorphic vision systems.
Figure 3: An overview of our framework. Our method consists of three parts: (a) Feature Initialization(Sec III-A), (b) Cross-modal Feature Restoration(Sec III-B), and (c) Frequency-aware Feature Fusion(Sec III-C). Specifically, Cross-modal Feature Restoration contains multiple Cross-RWKV blocks for feature alignment, and Frequency-aware Feature Fusion integrates image and event features for final output.
III Methodology
We propose EvRWKV, a novel framework designed for continuous cross-modal interaction through dual-domain processin that leverages the complementary advantages of RGB data and event streams to achieve robust illumination enhancement and noise suppression. As illustrated in Fig. 3, our framework comprises three key parts: 1) Feature Initialization, 2) Cross-modal Feature Restoration, 3) Frequency-aware Feature Fusion. The framework takes a low-light image $I$ and paired event stream $\{e_{k}\}_{k=1}^{N}$ as inputs, producing an enhanced image $I_{en}$. At first, a series of convolution layers extract features from low light images and event data, and then process them through the 4-level U-shaped encoder-decoder architecture composed of cross-RWKV composed of spatial and channel hybrid components. The output of the architecture passes the EISFE module is enhanced in frequency domain and spatial domain fusion processing to improve image quality.
III-A Feature Initialization
Recent LLIE methods demonstrate that preliminary enhancement facilitates subsequent restoration. Following Retinex theory [24], an image $I\in\mathbb{R}^{H\times W\times 3}$ can be decomposed into reflectance $R\in\mathbb{R}^{H\times W\times 3}$ and illumination $L\in\mathbb{R}^{H\times W}$ as $I=R\odot L$, where $\odot$ denotes element-wise multiplication. Following EvLight [30], we enhance the low-light image $I$ through:
$I_{lu}=I\odot\hat{L}+I,$ (1)
where $\hat{L}$ is the estimated illumination map, implementing rough enhancement while preserving original details. The event stream $\{e_{k}\}$ is converted into a voxel grid $E\in\mathbb{R}^{H\times W\times B}$ ($B=32$) to preserve spatio-temporal information. The image $I_{lu}$ and event voxel $E$ are both downsampled to $\frac{H}{2}\times\frac{W}{2}$ via a two-layer $\times3$ conv with strides 1 & 2, extracting features $X^{\text{img}},X^{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ .
III-B Cross-modal Feature Restoration
Existing methods for event-image fusion in low-light scenarios often fail to balance computational efficiency and robust cross-modal interaction, particularly in capturing high-frequency motion cues from events and preserving spatial details from images [17, 30]. Furthermore, they frequently treat event and image data as independent streams, failing to fully exploit the fine-grained spatiotemporal relationships between them. To address this, we propose the Cross-RWKV Module, which integrates Re-WKV Attention for capturing global interactions and Cross-Modal Switch Omni-Shift (CS-Shift) for local feature enhancement and multi-modal information interaction. Specifically, multiple Cross-RWKV Modules form a 4-level encoder-decoder. The encoder progressively downsamples spatial dimensions and expands channels, while the decoder upsamples to restore resolution, enabling multi-scale cross-modal fusion. The outputs $O_{\text{img}},O_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 2C}$ is split by a 3×3 convolution into $\hat{X}_{\text{img}},\hat{X}_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$. As shown in Fig.4, the module operates through spatial and channel mix, enabling efficient cross-modal fusion of event and image features. The details of the Cross-RWKV Block are as follows:
III-B1 Spatial Mix
The overall structure of Spatial Mix is illustrated in the Fig.4. This module aims to establish long-range dependencies across spatial dimensions while enabling effective interaction and information fusion between image and event modalities. Given the input feature sequences $X^{\text{img}}$, $X^{\text{ev}}\in\mathbb{R}^{T\times C}$, where $T=\frac{H}{2}\times\frac{W}{2}$ represents the spatial resolution of tokens, the module processes image tokens and event tokens in parallel, first performing spatial-level feature fusion through Layer Normalization (LN) and the CS-shift mechanism.
$X_{\text{s}}^{\text{img}},X_{\text{s}}^{\text{ev}}=\text{CS-Shift}(LN_{1}(X^{\text{img}}),LN_{2}(X^{\text{ev}})),$ (2)
where CS-shift is formulated as:
[…]
where $w_{i}^{img}$ and $w_{i}^{ev}$ represent learnable parameters for scaling specific branches, and $\text{Conv}_{k\times k}()$ denotes depthwise convolution with kernel size $k$. This CS-shift mechanism captures both local and global features by fusing information under different receptive fields, resulting in accurate token shift outcomes.
Following the shift operation, instead of using standard linear projections, we adopt a depthwise separable convolution to compute keys:
$K_{s}^{\text{img}}=\text{DPConv}(X_{s}^{\text{img}}),K_{s}^{\text{ev}}=\text{DPConv}(X_{s}^{\text{ev}}),$ (4)
where $\text{DPConv}()$ consists of Depthwise Convolution and Pointwise Convolution, designed to preserve fine-grained spatial structures while efficiently expanding the receptive field.
Meanwhile, the value and receptance vectors for each modality are generated through linear projections:
$V_{s}^{\text{img}}=X_{s}^{\text{img}}W_{v}^{\text{img}},\quad R_{s}^{\text{img}}=X_{s}^{\text{img}}W_{r}^{\text{img}},$ (5)
$V_{s}^{\text{ev}}=X_{s}^{\text{ev}}W_{v}^{\text{ev}},\quad R_{s}^{\text{ev}}=X_{s}^{\text{ev}}W_{r}^{\text{ev}},$ (6)
where $W_{\text{img}}^{v}$, $W_{\text{img}}^{r}$, $W_{\text{ev}}^{v}$, and $W_{\text{ev}}^{r}$ are four fully connected layers.
Following the Restore-RWKV [57], we employ the Re-WKV mechanism to compute the global attention result $\text{wkv}\in\mathbb{R}^{T\times C}$. At the same time, key representations are directly exchanged between image tokens and event tokens, allowing each modality to attend to the spatial structure of the other modality, thereby enabling joint spatial reasoning between images and events:
$\text{wkv}^{\text{img}}=\text{Re-WKV}(K_{s}^{\text{ev}},V_{s}^{\text{img}}),$ (7)
$\text{wkv}^{\text{ev}}=\text{Re-WKV}(K_{s}^{\text{img}},V_{s}^{\text{ev}})$ (8)
where Re-WKV is implemented by applying the Bi-WKV attention mechanism recurrently along both horizontal and vertical directions. Specifically, the attention output for the $t$-th token is computed as:
$\text{wkv}_{t}=\frac{\sum_{i=1,i\neq t}^{T}e^{-(|t-i|-1)/T\cdot(w+k_{i})}v_{i}+e^{u+k_{t}}v_{t}}{\sum_{i=1,i\neq t}^{T}e^{-(|t-i|-1)/T\cdot(w+k_{i})}+e^{u+k_{t}}}$ (9)
Figure 4: Architecture of the Cross-RWKV module, which includes Spatial Mix for spatial feature processing and Channel Mix for channel-wise interaction.
Here, $T$ denotes the total number of tokens, $k_{i}$ and $v_{i}$ are the key and value vectors at position $i$, and $w\in\mathbb{R}^{C}$, $u\in\mathbb{R}^{C}$ are learnable parameters that control the relative positional bias and provide additional weight to the current token, respectively. By applying Bi-WKV iteratively across multiple directions, Re-WKV effectively models long-range dependencies in 2D spatial structures, enabling enhanced global contextual reasoning across modalities.
To further promote cross-modal feature alignment, we introduce learnable fusion gates $\alpha_{\text{img}},\alpha_{\text{evt}}\in\mathbb{R}^{C}$, applied to modulate the contribution from each modality during the final representation fusion:
$X_{1}=\sigma(\alpha_{\text{img}})\cdot\text{wkv}^{\text{img}}+\left(1-\sigma(\alpha_{\text{img}})\right)\cdot X^{\text{ev}},$ (10)
$X_{2}=\sigma(\alpha_{\text{evt}})\cdot\text{wkv}^{\text{ev}}+\left(1-\sigma(\alpha_{\text{evt}})\right)\cdot X^{\text{img}}$ (11)
These fused representations are then gated by the learned receptance signals to produce the final outputs:
$O_{\text{img}}=\left(\sigma(R_{\text{img}})\odot X_{1}\right)W_{o}^{\text{img}},$ (12)
$O_{\text{evt}}=\left(\sigma(R_{\text{evt}})\odot X_{2}\right)W_{o}^{\text{evt}}$ (13)
W where $W_{o}^{\text{img}}$ and $W_{o}^{\text{evt}}$ are linear output projection matrices.
This design enables explicit cross-modal interaction through key exchange, spatially-aware recurrent updates via Re-WKV, and adaptive fusion through gating, yielding a robust mechanism for unified spatial modeling over both image and event streams.
Figure 5: Overview of the proposed EISFE module. (a) The overall structure includes Channel and Spatial Attention Blocks combined with an Adaptive Gaussian Filter to fuse inputs from image and event features. (b) Details of the Channel Attention Block and Spatial Attention Block architectures.
III-B2 Channel Mix
The Channel Mix module is designed to model cross-channel interactions within each spatial token while preserving the spatial structure of visual inputs. Unlike the Spatial Mix module that jointly processes both image and event modalities, Channel Mix focuses solely on the image stream. This design choice stems from the observation that channel-wise semantic structures are generally more informative and stable in image data than in sparse and noisy event representations.
Given the image features $X_{\text{img}}\in\mathbb{R}^{T\times C}$, where $T=H\times W$ denotes the number of spatial tokens and $C$ is the channel dimension, the input first undergoes an Omni-Shift operation to enhance local contextual information:
$X_{\text{c}}^{\text{img}}=\text{OmniShift}(\text{Conv}(X_{\text{img}}))$ (14)
The shifted feature is then passed through a gated channel-wise mixing pipeline. A key projection is first applied:
$K_{\text{c}}=\text{squared ReLU}(X_{\text{c}}^{\text{img}}W_{k})$ (15)
where $W_{k}\in\mathbb{R}^{C\times C_{h}}$ is the key projection matrix and the squared ReLU non-linearity emphasizes strong activations while suppressing weak ones. This is followed by a value projection to obtain the transformed feature:
$V_{\text{c}}=K_{\text{c}}W_{v},\quad W_{v}\in\mathbb{R}^{C_{h}\times C}$ (16)
A separate gating signal is generated through a linear transformation and a sigmoid activation:
$R_{\text{c}}=\sigma(X_{\text{c}}^{\text{img}}W_{r}),\quad W_{r}\in\mathbb{R}^{C\times C}$ (17)
The final output is obtained via element-wise modulation:
$O_{\text{c}}=R_{\text{c}}\odot V_{\text{c}}$ (18)
By restricting Channel Mix to the image modality only, we ensure more stable and informative feature transformations. Event streams, while rich in motion cues, often suffer from high sparsity and noise in the channel dimension, making them less suitable for dense feedforward modeling. Therefore, reserving event data for interaction in the spatial domain (via Spatial Mix) while using only image data for channel-wise transformations leads to a better division of modeling responsibilities, improving overall robustness and representation quality.
III-C Frequency-aware Feature Fusion
TABLE I: Quantitative comparison of different methods on two real captured event datasets, SDE dataset and RELED dataset. The best and second best results are highlighted in bold and underlined respectively.
Method Publication Backbone Inputs SDE-in SDE-out RELED
[…]
SNR-Net[55] CVPR2022 Transformer $\checkmark$ $\times$ 20.05 0.630 22.18 0.661 26.47 0.851
Uformer[51] CVPR2022 Transformer $\checkmark$ $\times$ 21.09 0.752 22.32 0.725 28.86 0.832
LLFlow-L-SKF[53] CVPR2023 CNN $\checkmark$ $\times$ 20.92 0.661 21.68 0.647 28.62 0.862
Retinexformer[3] ICCV2023 Transformer $\checkmark$ $\times$ 21.3 0.692 22.92 0.683 26.66 0.865
RetinexMac[33] TCSVT2025 Transformer $\checkmark$ $\times$ 20.61 0.650 21.68 0.683 28.55 0.852
eSL-Net[46] ECCV2020 CNN $\checkmark$ $\checkmark$ 21.42 0.725 21.39 0.681 25.76 0.796
ELIE[17] TMM2023 Transformer $\checkmark$ $\checkmark$ 19.98 0.617 20.69 0.653 26.62 0.862
LLVE-SEG[35] AAAI2023 Transformer $\checkmark$ $\checkmark$ 21.79 0.705 22.35 0.690 29.19 0.875
ELEDNet[20] ECCV2024 Transformer $\checkmark$ $\checkmark$ 21.46 0.7126 22.91 0.726 31.96 0.910
EvLight[30] CVPR2024 Transformer $\checkmark$ $\checkmark$ 22.21 0.758 22.13 0.725 31.29 0.865
Ours - RWKV $\checkmark$ $\checkmark$ 23.09 0.766 23.43 0.767 32.18 0.910
To effectively integrate complementary cues from image and event modalities, we propose the EISFE module, a Frequency-aware Image-Event Spatial-Frequency Enhancement module. This dual-domain fusion mechanism jointly exploits frequency-domain smoothness and spatial-domain adaptivity to enhance reconstruction quality, particularly in challenging scenarios such as motion blur, fast motion, or low-light degradation.
Given the restored image feature $\hat{X}_{\text{img}}$, event feature $\hat{X}_{\text{ev}}$, and raw image input $X^{\text{img}}$, we first concatenate them along the channel dimension to form a unified representation:
$X_{\text{fused}}=\text{Concat}(\hat{X}_{\text{img}},\hat{X}_{\text{ev}},X^{\text{img}}).$ (19)
We then apply two separate $1\times 1$ convolutions to decompose $X_{\text{fused}}$ into two parallel branches: a frequency branch $X_{\text{freq}}$ and a spatial branch $X_{\text{spat}}$.
In the frequency branch, we introduce a channel-wise Adaptive Gaussian Filtering mechanism to selectively suppress noise while preserving structural information. Unlike traditional fixed-filter approaches, our method learns a distinct Gaussian standard deviation $\sigma_{c}$ for each channel, enabling content-adaptive filtering across different frequency sensitivities. Given the frequency-branch input $X_{\text{freq}}\in\mathbb{R}^{C\times H\times W}$, we generate a 2D Gaussian kernel $G_{\sigma_{c}}\in\mathbb{R}^{K\times K}$ for each channel using its learned $\sigma_{c}\in[\sigma_{\min},\sigma_{\max}]$, and define it as:
$G_{\sigma_{c}}(x,y)=\frac{1}{2\pi\sigma_{c}^{2}}\exp\left(-\frac{x^{2}+y^{2}}{2\sigma_{c}^{2}}\right),$ (20)
where $(x,y)$ are spatial coordinates.
To efficiently apply filtering, we perform the convolution in the frequency domain via Fast Fourier Transform (FFT). Specifically, for each channel $c$, the filtered output is computed as:
$\hat{X}_{\text{freq},c}=\mathcal{F}^{-1}\left(\mathcal{F}(X_{\text{freq},c})\cdot\mathcal{F}(G_{\sigma_{c}})\right),$ (21)
where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote 2D FFT and inverse FFT, respectively. This formulation allows the model to perform global frequency-aware filtering with high efficiency and adaptivity.
By learning $\sigma_{c}$ per channel, this module dynamically adjusts filtering strength: it smooths flat areas while preserving edge structures and fine details, making it well-suited for denoising and contrast enhancement in degraded conditions.
In parallel, the spatial branch processes $X_{\text{spat}}$ using a deformable convolution. This allows flexible receptive fields that adapt to local content, enabling the model to capture fine-grained spatial variations such as motion boundaries or geometric distortions. The output is denoted as:
$\hat{X}_{\text{spat}}=\text{DeformConv}(X_{\text{spat}}),$ (22)
which retains high-frequency spatial patterns that complement the globally smoothed features from the frequency branch.
To integrate the outputs of the frequency and spatial branches, we design a hierarchical attention fusion strategy that sequentially applies spatial and channel attention mechanisms. This allows the model to selectively emphasize salient patterns while suppressing irrelevant noise, thereby enhancing cross-modal feature integration. As illustrated in Fig. 5, the fusion begins by computing spatial attention maps independently for both branches:
$X_{\text{attn-spat}}=A_{\text{freq}}\odot\hat{X}_{\text{freq}}+A_{\text{spat}}\odot\hat{X}_{\text{spat}},$ (23)
where $A_{\text{freq}},A_{\text{spat}}$ denotes spatial attention maps independently for both branches, and $\odot$ represents element-wise multiplication. This formulation adaptively combines spatial details and frequency-aware structures based on per-pixel relevance.
Following the spatial attention stage, we further refine the fused features through channel-wise attention to emphasize globally informative dimensions. As shown in Fig. 5, we first aggregate global context via adaptive global average pooling:
$X_{\text{attn}}=X_{\text{attn-spat}}\odot A_{\text{chan}}.$ (24)
This two-stage attention mechanism allows the EISFE module to dynamically adapt to both local and global feature importance. The spatial attention modulates each spatial location based on semantic content, while the channel attention globally reweighs feature maps to highlight informative channels. Together, these mechanisms enable robust and adaptive fusion of image-event representations under various degradation conditions.
Finally, the fused feature map $X_{\text{attn}}\in\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}$ is first linearly projected with a $1\times 1$ convolution to refine channel interactions, then up–sampled to the original spatial resolution by a learnable $4\times 4$ transposed convolution, and finally remapped to the RGB space through another $1\times 1$ convolution.
III-D Loss Function
We optimise EvRWKV in an end‑to‑end fashion by minimising a weighted combination of four complementary losses: the pixel‑wise reconstruction loss $\mathcal{L}_{r}$, the perceptual loss $\mathcal{L}_{p}$, the Structural Similarity loss $\mathcal{L}_{s}$, and its multi‑scale extension $\mathcal{L}_{m}$. Specifically, $\mathcal{L}_{\text{total}}$ is given by
$\mathcal{L}_{\text{total}}=\lambda_{r}\,\mathcal{L}_{r}+\lambda_{p}\,\mathcal{L}_{p}+\lambda_{s}\,\mathcal{L}_{s}+\lambda_{m}\,\mathcal{L}_{m},$ (25)
where the hyper‑parameters $\lambda_{r},\lambda_{p},\lambda_{s},\lambda_{m}$ balance the contributions of the individual terms.
To retain low‑level fidelity, we adopt the L1‑Charbonnier loss to penalise pixel errors between the enhanced image $I_{\mathrm{en}}$ and the ground‑truth image $I_{\mathrm{gt}}$:
$\mathcal{L}_{r}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(I_{\text{en}}-I_{\text{gt}})_{i}^{2}+\epsilon^{2}},$ (26)
where $I_{\mathrm{en}},I_{\mathrm{gt}}$ denote the enhanced and ground truth images and $\varepsilon=10^{-4}$ ensures numerical stability.
We employ the perceptual loss [25] to preserve the high-level similarity between the fused image and source images, which is given by
$\mathcal{L}_{p}=\bigl{\lVert}\,\Phi(I_{\mathrm{en}})-\Phi(I_{\mathrm{gt}})\,\bigr{\rVert}_{1},$ (27)
where $\Phi$ represents feature extraction using the Alex network [22].
For every $11\times 11$ Gaussian window $\Omega$ we compute the Structural Similarity (SSIM) index between the enhanced patch $x=I_{\mathrm{en}}\!\big{|}_{\Omega}$ and the corresponding ground-truth patch $y=I_{\mathrm{gt}}\!\big{|}_{\Omega}$:
$\operatorname{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (28)
where $\mu$, $\sigma$, and $\sigma_{xy}$ denote local means, standard deviations and cross-variance, while $C_{1}$ and $C_{2}$ avoid division by zero. Averaging the index over all windows and turning it into a loss gives
$\mathcal{L}_{s}=1-\operatorname{SSIM}(I_{\mathrm{en}},I_{\mathrm{gt}}).$ (29)
To capture structures at multiple resolutions, we build a five-level Gaussian pyramid and evaluate SSIM at every scale $j$. The resulting multi-scale similarity is the weighted geometric mean
$\operatorname{MS\text{-}SSIM}(x,y)=\prod_{j=1}^{5}\operatorname{SSIM}_{j}(x,y)^{\,\omega_{j}},$ (30)
where $\{\omega_{j}\}$ are the standard MS-SSIM weights ($\sum_{j}\omega_{j}=1$). The corresponding loss is obtained by
$\mathcal{L}_{m}=1-\operatorname{MS\text{-}SSIM}(I_{\mathrm{en}},I_{\mathrm{gt}}).$ (31)
IV EXPERIMENTS
IV-A Experimental Settings
IV-A1 Datasets and Evaluation Metrics
SED [30], SDSD [47], and RELED [20] datasets are selected to test the enhancement performance of our EvRWKV framework.
SED dataset contains 91 image–event paired sequences (43 indoor and 48 outdoor) captured with a DAVIS346 event camera at a resolution of $346\times 260$; it is primarily used to evaluate the capability of different methods to leverage event cues for LLIE under diverse illumination conditions.
SDSD dataset provides 150 low-light/normal-light paired video sequences with an original resolution of $1920\times 1080$; following the dynamic-subset split (125 sequences for training and 25 for testing), all videos are down-sampled to $346\times 260$, and their event streams are synthesized with the v2e [15] simulator. This dataset evaluates the robustness of enhancement methods in dynamic low-light scenes.
RELED dataset comprises 42 scenes (29 for training and 13 for testing), each offering a low-light blurred input directly paired with a sharp, normal-light reference, making it suitable for benchmarking algorithms that jointly address illumination enhancement and deblurring.
Figure 6: Visual comparisons with state-of-the-art methods on the SDE dataset.
Figure 7: Visual comparisons with state-of-the-art methods on the SDSD dataset.
Together, these three benchmarks provide a comprehensive testbed for assessing the multi-modal LLIE capability of our EvRWKV method. We use the peak-signal-to-noise ratio (PSNR) [14] and SSIM [52] for evaluation.
TABLE II: Quantitative comparison of different methods on the SDSD dataset. The best and second best results are highlighted in bold and underlined respectively.
Method Publication SDSD-in SDSD-out
$uparrow PSNR\uparrow SSIM\uparrow$
SNR-Net[55] CVPR’22 24.74 0.830 24.82 0.740
Uformer[51] CVPR’22 24.03 0.900 24.08 0.818
LLFlow-L-SKF[53] CVPR’23 23.39 0.818 20.39 0.634
Retinexformer[3] ICCV’23 25.90 0.852 26.08 0.815
RetinexMac[33] TCSVT’25 27.61 0.900 24.75 0.783
ELIE[17] TMM’23 27.46 0.879 23.29 0.742
eSL-Net[46] ECCV’20 23.68 0.821 24.95 0.805
LLVE-SEG[35] AAAI’23 27.58 0.888 23.51 0.726
ELEDNet[20] ECCV’24 28.52 0.912 25.08 0.828
EvLight[30] CVPR’24 28.49 0.909 25.17 0.817
Ours - 28.96 0.920 27.93 0.839
IV-A2 Implementation Details
All experiments were conducted with the Adam optimizer [21]. Learning rates were set to 1e41e{-4} for SDE datasets, 1.5e41.5e{-4} for SDSD datasets, and 1e41e{-4} for RELED datasets. The framework was trained on an NVIDIA RTX 3090 GPU for 80 training cycles with a batch size of eight. For the training set, we employed data-augmentation techniques consisting of random $256\times 256$ crops, horizontal flips, and rotations of $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$. For the test set, we applied only a center crop of $256\times 256$.
IV-A3 State-of-the-Art Methods for Comparisons
We compared our EvRWKV with ten methods, including five methods that only use RGB images as input(SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], RetinexMac [33], Retinexformer [3]), five methods that use a RGB image and paired events as inputs(ELIE [17], eSL-Net [46], LLVE-SEG [35], ELEDNet [20], EvLight [30]). Since the source code of LLVE-SEG [35] is not publicly available, we implemented it according to the description in its paper. For the SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], RetinexMac [33], Retinexformer [3], ELIE [17], eSL-Net [46], ELEDNet [20] and EvLight [30] methonds, we used the codes released by the authors to output their results.
IV-B Quantitative and Qualitative Evaluation
We evaluate the proposed EvRWKV framework on three challenging datasets—SDE, SDSD, and RELED—covering diverse LLIE scenarios. SDE and RELED contain real-world event-image pairs captured under indoor/outdoor and low-light blurry conditions respectively, while SDSD consists of synthetic events generated from dynamic video sequences. We use PSNR and SSIM as quantitative metrics and provide detailed qualitative analysis to validate the effectiveness of our approach.
IV-B1 Evaluation on the SDE Dataset
For LLIE on the SDE dataset, the ability to simultaneously suppress noise and preserve fine structural details is the key evaluation criterion. We first qualitatively compare our EvRWKV method with several representative approaches including eSL-Net [46], ELEDNet [20], EvLight [30], and RetinexMac [33]. As shown in Fig. 6, RetinexMac and ELEDNet reduce noise to some extent but tend to oversmooth textures, resulting in loss of important details in dark regions. As shown in Fig. 6, EvLight maintains more structural information but suffers from residual noise and blur artifacts. eSL-Net, with a relatively small model size, provides limited enhancement capabilities and often underperforms in challenging scenes. In contrast, our EvRWKV method achieves a better balance between noise suppression and texture preservation, producing cleaner, brighter images with rich details and minimal artifacts. This demonstrates the effectiveness of our dual-domain fusion and adaptive gating mechanisms in handling complex low-light scenarios.
In terms of quantitative evaluation, Table I lists the PSNR and SSIM scores of different methods on the SDE dataset under both SDE-in and SDE-out settings. Our method achieves the highest scores of 23.09 PSNR and 0.766 SSIM for SDE-in, as well as 23.43 PSNR and 0.767 SSIM for SDE-out. These results outperform all compared methods including SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], LLVE-SEG [35], Retinexformer [3], ELIE [17], LLVE-SEG [35], and the aforementioned qualitative comparison models. The significant improvements in both perceptual quality and objective metrics validate the superiority of our EvRWKV framework in LLIE.
IV-B2 Evaluation on the SDSD Dataset
We further qualitatively compare the performance of different methods on the SDSD dataset for LLIE under challenging lighting and motion conditions. In Fig. 7, we observe that most methods improve the visibility and brightness of degraded images. However, RetinexMac [33], ELEDNet [20], and EvLight [30] exhibit limitations: As shown in Fig. 7, RetinexMac and ELEDNet tend to over-smooth textures and lose fine details, While EvLight, although preserving more structural information, causes noticeable color distortion in some regions. As shown in Fig. 7, eSL-Net [46] shows relatively limited enhancement capability due to its smaller model size. In comparison, our EvRWKV method delivers superior noise suppression and detail preservation, effectively enhancing both static and dynamic scene elements without introducing artifacts. The enhanced images show clearer textures, better contrast, and natural brightness, demonstrating the robustness of our dual-domain cross-modal fusion strategy.
Table II presents the PSNR and SSIM scores of different methods evaluated on the SDSD dataset. Our method achieves the highest scores in both SDSD-in and SDSD-out scenarios, with PSNR values of 28.96 and 27.93, and SSIM values of 0.920 and 0.839, respectively. These results outperform all other compared methods, validating the strong restoration ability and noise robustness of our EvRWKV framework under various low-light conditions. Overall, our approach demonstrates both qualitative and quantitative advantages over existing state-of-the-art techniques on the SDSD dataset.
IV-B3 Evaluation on the RELED Dataset
Figure 8: Visual comparisons with state-of-the-art methods on the Reled dataset.
TABLE III: Ablation study of key components on the SDSD-in dataset.
[…]
w/o EISFE 27.35 0.894
w/o SpatialMix 26.88 0.798
w/o ChannelMix 26.60 0.896
Ours 28.96 0.920
To comprehensively evaluate the enhancement performance of our EvRWKV method on the RELED dataset, we compare it qualitatively with several representative approaches including eSL-Net [46], ELEDNet [20], EvLight [30], and RetinexMac [33]. As shown in Fig. 8, the RELED dataset contains extremely low-light images with severe motion blur and noise. As shown in Fig. 8, eSL-Net often causes overexposure while attempting to enhance brightness. RetinexMac boosts overall brightness but fails to eliminate noticeable blur. ELEDNet and EvLight aim to balance illumination, but this often leads to the loss of finer details. In contrast, our EvRWKV method effectively leverages event-based motion cues and image textures to suppress noise, reduce blur, and restore fine details, producing visually clearer, sharper, and more natural images without color artifacts or overexposure.
Table I reports the PSNR and SSIM scores for different methods on the RELED dataset. Our EvRWKV achieves the best quantitative results with a PSNR of 32.18 and SSIM of 0.910, significantly outperforming other methods including eSL-Net, ELEDNet, EvLight, and RetinexMac. These quantitative gains demonstrate the robustness and generalization capability of our dual-domain cross-modal fusion framework in handling challenging low-light and motion-degraded scenarios. Overall, our method achieves excellent qualitative and quantitative results across all three real-world LLIE benchmarks.
IV-C Ablation Study
We conduct comprehensive ablation studies on the SDSD-in dataset to evaluate the contribution of each key component in our EvRWKV framework. In particular, we assess the effectiveness of the EISFE module, the SpatialMix and ChannelMix mechanisms within the Cross-RWKV module, and the role of each individual loss term in the overall objective function. For each study, we remove or modify a single module while keeping the rest of the framework unchanged to isolate its impact. The results clearly demonstrate the necessity and complementarity of these components in achieving robust and high-quality LLIE. As shown in Fig. 9, the removal of EISFE (b) results in color distortion. Without SpatialMix (c), visible noise and blurred textures appear, while omitting ChannelMix (d) leads to flat tonal rendering. The full model (e) demonstrates the best visual clarity and color fidelity, especially in complex shadow regions.
IV-C1 Impact of EISFE Module
To assess the contribution of frequency-aware fusion, we remove the EISFE module and directly decode the features after the Cross-RWKV backbone. As shown in Table III, the absence of EISFE leads to a 1.61 dB drop in PSNR and 0.026 reduction in SSIM. This indicates the importance of frequency-domain filtering and spatial deformable convolution in denoising and structure restoration. Visually, the model without EISFE shows residual noise and blurred edges.
IV-C2 Impact of SpatialMix and ChannelMix
To verify the effectiveness of the spatial and channel interaction designs in Cross-RWKV, we conduct ablation studies by individually removing each module. As shown in Table III, removing the SpatialMix results in a 2.08 dB PSNR drop and 0.122 SSIM decline, highlighting its critical role in capturing long-range dependencies and facilitating event-image alignment. On the other hand, removing ChannelMix leads to a 2.36 dB decrease in PSNR and a 0.024 SSIM reduction, indicating its importance in refining semantic consistency across feature channels. These results confirm that both modules contribute distinctly and significantly to performance. As illustrated in Fig. 9, ChannelMix aids in preserving local contrast, while SpatialMix enhances spatial structure and reduces texture inconsistency.
Figure 9: Visualization of ablation results.
IV-C3 Impact of Loss Functions
To evaluate the contribution of each loss term, we progressively remove pixel, perceptual, SSIM, and MS-SSIM losses from the training objective. The results are summarized in Table IV. Removing the reconstruction loss $L_{r}$ results in the lowest performance, with a PSNR of 26.84 and an SSIM of 0.896. Excluding the perceptual loss $L_{p}$ leads to a slight reduction in performance, with a PSNR of 26.18 and an SSIM of 0.885. Removing the SSIM loss $L_{s}$ results in a PSNR of 26.06 and an SSIM of 0.876, reflecting a loss in structural preservation. Finally, excluding the multi-scale SSIM loss $L_{m}$ leads to a PSNR of 27.84 and an SSIM of 0.899. The full model, with all losses included, achieves the best results with a PSNR of 28.96 and an SSIM of 0.920, demonstrating a 0.84 dB gain in PSNR over the setting without $L_{r}$. This confirms the complementary benefits of each loss term in guiding high-quality reconstruction.
TABLE IV: Ablation study of loss functions on the SDSD-in dataset.
[…]
w/o reconstruction loss ($L_{r}$) 26.84 0.896
w/o perceptual loss ($L_{p}$) 26.18 0.885
w/o SSIM loss ($L_{s}$) 26.06 0.876
w/o MS-SSIM loss ($L_{m}$) 27.84 0.899
Ours 28.96 0.920
IV-D Limitations
While EvRWKV performs well across diverse low-light conditions, it encounters challenges in extremely dark scenes with sparse event activity and limited structural cues. In such cases, the framework may struggle to recover fine details or accurate colors, occasionally leading to over-smoothing or artifacts. Additionally, severe motion blur or sensor noise can impair spatial-temporal alignment. Future work will focus on improving robustness through uncertainty modeling and adaptive fusion strategies for broader generalization.
V Conclusion
In this paper, we proposed EvRWKV, a novel and effective framework for LLIE that leverages continuous cross-modal interaction between event streams and low-light images through dual-domain processing. Our approach uniquely integrates the RWKV architecture with the EISFE module, enabling adaptive and robust fusion of asynchronous event data and standard images. The Cross-RWKV module effectively captures fine-grained temporal and spatial dependencies, while the EISFE module jointly performs noise suppression and structural restoration via frequency-domain filtering and spatial deformable convolutions. Extensive experiments on multiple challenging real-world low-light datasets demonstrate that EvRWKV significantly outperforms existing state-of-the-art methods in both quantitative metrics and visual quality, effectively suppressing noise, enhancing contrast, and restoring fine details even under severe degradation such as motion blur and extreme darkness. Our dual-domain fusion and adaptive gating mechanisms successfully balance complementary information from event and image modalities, resulting in superior enhancement robustness and generalization. In future work, we plan to explore further improvements in cross-modal fusion strategies and extend the application of the EvRWKV framework to other challenging imaging tasks such as real-time video enhancement and dynamic scene understanding in extreme lighting conditions.